# RAGDiffusion: Faithful Cloth Generation via External Knowledge Assimilation

### v1

@@ 6 Additional analysis: enhancing generalization through external retrieval databases
RAGDiffusion is trained on upper-body clothing data and has not encountered lower-body garments during training. In this testing phase, we collect embeddings and corresponding landmarks for $856$ lower-body items and incorporate them into the external memory database. Subsequently, we gather $50$ lower-body in-the-wild clothing samples as a test set, using a ReferenceNet version as a baseline for comparison. The results in Fig. 11 demonstrate that retrieval significantly improves generalization capabilities. By injecting the embeddings of lower-body garments along with their corresponding contour landmarks as conditional constraints, RAGDiffusion produces accurate results in the lower-body domain, showcasing its strong out-of-distribution (OOD) compatibility. In contrast, the ReferenceNet version is noticeably confused by the concept of lower-body garments and fails to yield meaningful garment structures. This boost in generalization increases the operational maturity of RAGDiffusion, enabling it to effectively handle various OOD images submitted by users.
@@ 7 Limitations and future work
We present RAGDiffusion, an efficient RAG framework that supports the generation of standardized clothing assets by effectively addressing the prevalent challenges of structural hallucinations and fidelity in generated images. However, there are several limitations to discuss, which will help us improve the proposed framework further.
@@
Firstly, RAGDiffusion does not utilize an end-to-end training approach. Instead, it comprises four distinct processes: constructing the embedding space through contrastive learning, adjusting the U-Net backbone based on PGEA, establishing the memory database, and training the diffusion model. These separate processes impose higher demands on researchers’ expertise and coding capabilities. However, this modularity is inherent to the use of retrieval-augmented techniques, which necessitate multiple stages. Maintaining independent memory databases, independent latent spaces, and separate diffusion training ensure that the memory database can be freely adjusted as needed, effectively decoupling it from the diffusion model.
Secondly, we have observed a possibility of color bias in nearly pure-colored garments, particularly under very bright or very dark situations. This phenomenon is commonly encountered in image-based image editing, as the reconstruction loss constraints of Stable Diffusion are not particularly sensitive to color discrepancies. Additionally, the illumination variance can further impact the perceived colors of clothing. We note that extended training durations, along with the incorporation of contrast-enhancing data augmentation techniques, can partially alleviate this issue.
@@ 9.1 StructureNet
As shown in Tab. 4, we employ the Qwen2-VL-7B [5] language model to extract a total of $14$ clothing attributes, of which $10$ contribute to the construction of the garment structure embeddings, indicated by an asterisk *. The language model also provides captions describing the clothing content, thereby enhancing our external knowledge. We utilize $4$ NVIDIA 80GB-A100 GPUs to deploy the Qwen2-VL-7B service, achieving the attribute and caption generation time of under 3.5 seconds.
@@
In the construction of StructureNet, we first extracte features $f_{img}$ from in-the-wild clothing images or flat-lay images using the CLIP-ViT-L/14 [50] image encoder backbone (specifically from the second-to-last layer), resulting in a feature dimension of 768. Subsequently, we assign a 32-dimensional learnable embedding to each attribute extracted from the language model and concatenate the attribute features $f_{attr}$ with the image features $f_{img}$ to form a vector in $\mathbb{R}^{768+10\times 32}$. Through the Resampler module [69], we conduct a nonlinear mapping, ultimately forming an embedding $e$. The Resampler module consists of 4 layers of MLP and 4 layers of attention mechanisms, derived from the IP-Adapter open-source code. For features originating from different image domains $(x_{itw},x_{std})$, we designate the extracted embeddings as $(e_{itw},e_{std})$. The dual-tower ViT image encoder combined with the Resampler module is referred to as StructureNet. As described in the main body of paper, StructureNet is trained on STGarment using contrastive learning for 4 days on 4 NVIDIA 80GB-A100 GPUs with a batch size of 128.
@@ 9.2 EP-Adapter
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})+\mathsf{attention}(Q,K_{\text{image}},V_{\text{image}}).$
@@ 9.3 Guideline for collecting retrieval database
To construct a memory database for retrieval, we base our selection on the training set of 65,131 standard flat-lay garments. Initially, we filter approximately 15,000 samples from the original dataset according to the principle of category balance. Subsequently, we encode the samples into embeddings and employ Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [ester1996density] to cluster the samples. This step serves to eliminate potential noisy and non-standard data from the original dataset, thereby mitigating the adverse effects that outliers may have on retrieval results. We then perform random downsampling on the densely distributed samples to filter out overly similar instances. Through these operations, we aim to establish a high-quality, broadly representative standard flat-lay (embedding, garment) retrieval database with relatively low redundancy. By adjusting different clustering and downsampling parameters, we can generate retrieval libraries of four different scales: 1,000, 2,000, 4,000, and 8,000 samples, which are utilized in Section 4.2 of the main body of the paper.
@@ 10 Preliminary
The Image Prompt Adapter (IP-Adapter) is utilized to condition the Text-to-Image (T2I) diffusion model with a reference image for style control or content indication. This is typically achieved through global control at a high-level semantic layer. Specifically, it extracts features using an image encoder (e.g., the CLIP [50] image encoder) and incorporates an additional cross-attention layer onto the invariant text conditioning. Here, we denote $Q\in\mathbb{R}^{N\times d}$ as the query matrices extracted from the intermediate features of the U-Net, while $K_{\text{text}}\in\mathbb{R}^{N\times d}$ and $V_{\text{text}}\in\mathbb{R}^{N\times d}$ represent the key and value matrices derived from the prompt embeddings, where $N$ signifies the batch size. The cross-attention layer of the text branch is computed as follows:
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})=\mathsf{softmax}\left(\frac{QK_{\text{text}}^{\top}}{\sqrt{d}}\right).$
@@
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})+\mathsf{attention}(Q,K_{\text{image}},V_{\text{image}}).$
During training, the weights of the original U-Net are frozen, and only the projection layers of the key and value matrices in the image encoding branch, as well as the linear projection layer that maps the CLIP image embeddings, can be updated.
@@ 11 Experiment
11.1 Alignment during evaluation protocols.
Considering that our task can essentially be viewed as reconstructing a standard flat lay from a conditioning image, we have observed that the values of SSIM and LPIPS for this task are significantly worse than those typically seen in common reconstruction tasks. Therefore, we investigate the underlying reasons affecting the performance of SSIM and LPIPS. As shown in Fig. 16, we find that (1) the positioning of the generated flat-lay garment may influence the results of SSIM and LPIPS. Due to the lack of explicit positioning, the generated garment may deviate from the ground truth cloth in both horizontal and vertical directions, or exhibit minor differences in scale, which adversely impacts SSIM and LPIPS. (2) Even when dealing with a standard flat-lay garment, there may still be some discrepancies in the state of the garment. Factors such as the extent to which the sleeves are spread, the shooting angle, and lighting conditions can significantly affect the measurements of SSIM and LPIPS, leading to worse values for these metrics in the context of standard garment generation tasks compared to conventional reconstruction tasks.
To address the reconstruction errors caused by misalignment (Reason 1), we systematically crop and align the generated images with the ground truth (GT) images according to the bounding boxes, as illustrated in the third column of the Fig. 16. The data presented in the tables of the main body of the paper have all undergone this alignment process.
@@ 11.3 More visual results
We provide more visual results for inspection in Fig. 17 to demonstrate that RAGDiffusion synthesizes structurally and detail-faithful clothing assets. Fig. 14 provides more results about ablation study.

### v2

@@ 6.1 Enhanced generalizability due to RAG
RAGDiffusion is trained on upper-body clothing data and has not encountered lower-body garments during training. In this testing phase, we collect embeddings and corresponding landmarks for $856$ lower-body/dress items and incorporate them into the external memory database. Subsequently, we gather $50$ lower-body/dress in-the-wild clothing samples as a test set, using a ReferenceNet version as a baseline for comparison. The results in Fig. 10 demonstrate that retrieval significantly improves generalization capabilities.
By injecting the embeddings of lower-body garments along with their corresponding contour landmarks as conditional constraints, RAGDiffusion produces accurate results in the lower-body domain, showcasing its strong out-of-distribution (OOD) compatibility. In contrast, the ReferenceNet version is noticeably confused by the concept of lower-body garments and fails to yield meaningful garment structures. This is because current generative models necessitate corresponding training data to perform well. Other baselines face similar issues. ReferenceNet requires retraining with lower-body try-on training data to work effectively on lower garments. In contrast, RAGDiffusion does not need retraining or expensive try-on training data, as it can achieve results with just standard clothing database update, illustrating its greater flexibility. Of course, RAGDiffusion could perform better on lower garments if it underwent training similar to that on upper garments. This boost in generalization increases the operational maturity of RAGDiffusion, enabling it to effectively handle various OOD images submitted by users.
Generalization on Extreme Clothing Types. As RAGDiffusion works with real-world data, we provide more results to showcase how it handles super varied garments—like crazy patterns or funky designs in Fig. 9. I. Crazy patterns. Due to the pattern-level and detail-level faithful generation pipeline (Section 3), RAGDiffusion delivers accurate texture and logos even if the garment has crazy patterns, as shown in Main. Fig. 5, 6, and Fig. 9. II. Funky shapes. Highly varied shapes pose a significant challenge for all methods. Existing methods are nearly impossible to succeed in these scenarios. While our model still has a probability of generating distorted shapes, it also produces correct results sometimes, and overall, RAGDiffusion significantly outperforms the baselines.
6.2 Human-interpretable control through landmark manipulation due to RAG
In practice, the retrieval-acquired silhouette mask provides users with a visual shape preview opportunity before UNet denoising. Users can modify the final visual presentation by replacing recommended silhouette masks/landmarks (i.e., selecting from the recommended $K$ nearest neighbor landmark candidates during retrieval) before UNet denoising. For instance, in formal shirt scenarios, complete flattening may be required to convey a serious aesthetic, whereas in hoodie cases, users might prefer slightly bent sleeves with mild surface wrinkles to achieve a casual and relaxed appearance. As shown in  Fig. 11, we demonstrate several style modification cases of flat-lay garments through human-interpretable manipulation. This functionality is absent in end-to-end generative models like ReferenceNet [23], TryOffDiff [51], and TryOffAnyone [53], highlighting RAGDiffusion’s advantages in human-interpretable control and retrieval-based recommendation manipulation.
@@ 7 Limitations and future work
We present RAGDiffusion, an efficient RAG framework that supports the generation of standardized clothing assets by effectively addressing the prevalent challenges of structural hallucinations and fidelity in generated images. However, there is also limitations to discuss, which will help us improve the proposed framework further.
@@
Actually, we have observed a possibility of color bias in nearly pure-colored garments, particularly under very bright or very dark situations, as shown in Fig. 12. This phenomenon is commonly encountered in image-based image editing, as the reconstruction loss constraints of Stable Diffusion are not particularly sensitive to color discrepancies. Additionally, the illumination variance can further impact the perceived colors of clothing. We note that extended training durations, along with the incorporation of contrast-enhancing data augmentation techniques, can partially alleviate this issue.
@@ 8 Downstream applications of the standard garments
Retrieval Time Denoising Time LLM Caption Time
0.3 seconds 8.0 seconds 3.5 seconds
Table 5: Sampling time cost on an RTX 3090 GPU at resolution of $768\times 768$. Retrieval cost is a negligible part of generation.
@@ 9 More implementation details
9.1 Computational costs report
The computational overhead is relatively negligible when facing a larger retrieval memory database. In practical generation, we store feature vectors of a standard garment set locally as the retrieval memory database. During sampling, we retrieve the most similar garment vectors efficiently with FAISS [15] from the memory database in parallel based on the feature vector of the input image. On the other hand, the speed bottleneck of the generation process lies in the iterative denoising process, where retrieval time is a negligible part, as shown in Tab. 5.
@@ 9.2 StructureNet
As shown in Tab. 6, we employ the Qwen2-VL-7B [2] language model to extract a total of $14$ clothing attributes, of which $10$ contribute to the construction of the garment structure embeddings, indicated by an asterisk *. The language model also provides captions describing the clothing content, thereby enhancing our external knowledge. We utilize 44 NVIDIA H20 GPUs to deploy the Qwen2-VL-7B service, achieving the attribute and caption generation time of under 2 seconds.
@@
In the construction of StructureNet, we first extract features $f_{img}$ from in-the-wild clothing images or flat-lay images using the CLIP-ViT-L/14 [39] image encoder backbone (specifically from the second-to-last layer), resulting in a feature dimension of 768. Subsequently, we assign a 32-dimensional learnable embedding to each attribute extracted from the language model and concatenate the attribute features $f_{attr}$ with the image features $f_{img}$ to form a vector in $\mathbb{R}^{768+10\times 32}$. Through the Resampler module [57], we conduct a nonlinear mapping, ultimately forming an embedding $e$. The Resampler module consists of 4 layers of MLP and 4 layers of attention mechanisms, derived from the IP-Adapter open-source code. For features originating from different image domains $(x_{itw},x_{std})$, we designate the extracted embeddings as $(e_{itw},e_{std})$. The dual-tower ViT image encoder combined with the Resampler module is referred to as StructureNet. As described in the main body of paper, StructureNet is trained on STGarment using contrastive learning for 4 days on 4 NVIDIA H20 GPUs with a batch size of 128.
@@ 9.3 EP-Adapter
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})+\mathsf{attention}(Q,K_{\text{image}},V_{\text{image}}).$ (5)
@@ 9.4 Guideline for collecting retrieval database
To construct a memory database for retrieval, we base our selection on the training set of 65,131 standard flat-lay garments. Initially, we filter approximately 15,000 samples from the original dataset according to the principle of category balance. Subsequently, we encode the samples into embeddings and employ Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [16] to cluster the samples. This step serves to eliminate potential noisy and non-standard data from the original dataset, thereby mitigating the adverse effects that outliers may have on retrieval results. We then perform random downsampling on the densely distributed samples to filter out overly similar instances. Through these operations, we aim to establish a high-quality, broadly representative standard flat-lay (embedding, garment) retrieval database with relatively low redundancy. By adjusting different clustering and downsampling parameters, we can generate retrieval libraries of four different scales: 1,000, 2,000, 4,000, and 8,000 samples, which are utilized in Section 4.2 of the main body of the paper.
@@ 10 Preliminary
The Image Prompt Adapter (IP-Adapter) is utilized to condition the Text-to-Image (T2I) diffusion model with a reference image for style control or content indication. This is typically achieved through global control at a high-level semantic layer. Specifically, it extracts features using an image encoder (e.g., the CLIP [39] image encoder) and incorporates an additional cross-attention layer onto the invariant text conditioning. Here, we denote $Q\in\mathbb{R}^{N\times d}$ as the query matrices extracted from the intermediate features of the UNet, while $K_{\text{text}}\in\mathbb{R}^{N\times d}$ and $V_{\text{text}}\in\mathbb{R}^{N\times d}$ represent the key and value matrices derived from the prompt embeddings, where NN signifies the batch size. The cross-attention layer of the text branch is computed as follows:
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})=\mathsf{softmax}\left(\frac{QK_{\text{text}}^{\top}}{\sqrt{d}}\right)\cdot V_{\text{text}}.$ (7)
@@
$\mathsf{attention}(Q,K_{\text{text}},V_{\text{text}})+\mathsf{attention}(Q,K_{\text{image}},V_{\text{image}}).$ (8)
During training, the weights of the original UNet are frozen, and only the projection layers of the key and value matrices in the image encoding branch, as well as the linear projection layer that maps the CLIP image embeddings, can be updated.
@@ 11 Experiment
11.1 Alignment during evaluation protocols
Considering that our task can essentially be viewed as reconstructing a standard flat lay from a conditioning image, we investigate the potential reasons affecting the performance of SSIM and LPIPS. As shown in Fig. 17, we find that (1) the positioning of the generated flat-lay garment may influence the results of SSIM and LPIPS. Due to the lack of explicit positioning, the generated garment may deviate from the ground truth cloth in both horizontal and vertical directions, or exhibit minor differences in scale, which impacts SSIM and LPIPS. (2) Even when dealing with a standard flat-lay garment, there may still be some discrepancies in the state of the garment. Factors such as the extent to which the sleeves are spread, the shooting angle, and lighting conditions can affect the measurements of SSIM and LPIPS.
To address the measurement errors caused by misalignment (Reason 1), we systematically crop and align the generated images with the ground truth (GT) images according to the bounding boxes, as illustrated in the third column of the Fig. 17. The data presented in the tables of the main body of the paper have all undergone this alignment process.
11.2 About ControlNet baseline
We set the in-the-wild images as the input to the ControlNet conditioning branch, due to the structural aligned mask/canny images are absent without our proposed RAG. Actually, a straightforward ControlNet cannot handle this task well, as the input in-the-wild images do not have structural alignment with the desired output images.
@@ 11.4 More visual results
Fig. 15 provides more results about ablation study.
Fig. 18 provides more results on STGarment for inspection to demonstrate that RAGDiffusion synthesizes structurally and detail-faithful clothing assets.
Fig. 19 provides more cross dataset visual results on the unseen dataset Viton-HD, DressCode and the untrained categories lower-body/dresses from RAGDiffusion to validate the enhanced generalizability due to RAG.
@@
Figure 19: More cross dataset visual results on the unseen dataset Viton-HD, DressCode and the untrained categories lower-body/dresses from RAGDiffusion to validate the enhanced generalizability due to RAG. Best viewed when zoomed in.